### v2

# LADY: Linear Attention for Autonomous Driving Efficiency without Transformers

Abstract
End‑to‑end paradigms have demonstrated great potential for autonomous driving. Additionally, most existing methods are built upon Transformer architectures. However, transformers incur a quadratic attention cost, limiting their ability to model long spatial and temporal sequences—particularly on resource‑constrained edge platforms. As autonomous driving inherently demands efficient temporal modeling, this challenge severely limits the deployment and their real-time performance. Recently, linear attention mechanisms have gained increasing attention due to their superior spatiotemporal complexity. However, existing linear attention architectures are limited to self-attention, lacking support for cross-modal and cross-temporal interactions—both crucial for autonomous driving. In this work, we propose LADY, the first fully linear attention-based generative model for end-to-end autonomous driving. LADY enables fusion of long-range temporal context at inference with constant computational and memory costs, regardless of the history length of camera and LiDAR features. Additionally, we introduce a lightweight linear cross-attention mechanism that enables effective cross-modal information exchange. Experiments on the NAVSIM and Bench2Drive benchmarks demonstrate that LADY achieves state-of-the-art performance with constant-time and memory complexity, offering improved planning performance and significantly reduced computational cost. Additionally, the model has been deployed and validated on edge devices, demonstrating its practicality in resource-limited scenarios.
I Introduction
With the rapid advances in artificial intelligence, autonomous driving has gained widespread attention from both academia and industry, promising safer roads and more efficient transportation [zhao2024auto, li2024end, zheng2025genad, yu2025e2e]. While localization and scene perception have seen impressive progress [zhao2024survey, lu2022real, wang2023unitr, tian2024drivevlm, chen2024video], we focus on the critical challenge of motion planning [pan2024safe, wang2024llm], whose robustness directly determines ride safety. The mainly task of motion planning is to generate a collision-free, smooth, and dynamically feasible trajectory in real time [zhao2024survey]. Traditional rule-based planners rely on hand-crafted heuristics and careful tuning for corner cases, making them brittle when faced with the organic complexity of real-world traffic [teng2023motion]. In contrast, end-to-end learning-based planners map raw sensor inputs directly to feasible trajectories, offering greater scalability and robustness [chib2023recent, li2024ego].
Most existing end-to-end methods—such as Transfuser [chitta2022transfuser], UniAD [hu2023planning], FusionAD [ye2023fusionad] and VAD [jiang2023vad] imitate expert drivers to generate a single “best” trajectory, as shown in Fig. 1(a). However, human driving behavior is inherently multi-modal: in any given scenario, multiple plausible maneuvers may exist (e.g., yield vs. overtake). Therefore, some researchers propose generating multi-modal trajectory candidates to better capture diverse driving behaviors, as shown in Fig. 1(a). VADv2 [chen2024vadv2] and the Hydra-MDP series [li2024hydra, li2025hydra] introduce a large, fixed vocabulary of anchor trajectories and sample from these anchors to generate multiple feasible driving maneuvers. Inspired by the diffusion policy [chi2023diffusion], DiffusionDrive [liao2024diffusiondrive] introduces the diffusion process to the field of end-to-end autonomous driving and further adopts a truncated diffusion process to mitigate mode collapse and reduce computational overhead, yielding a set of varied, high-quality multi-modal feasible candidate trajectories.
Figure 1: Comparison of different end-to-end paradigms: (a - b) Existing approaches typically rely on transformers for attention-especially cross-attention versus (c) our approach is built entirely on linear attention.
However, existing methods typically rely on transformer for both feature fusion and cross-attention, imposing significant computational overhead for real-time applications, as shown in Fig. 1(a). Recently, DRAMA [yuan2024drama] proposes a Mamba-based encoder-decoder for efficient fusion of camera and LiDAR features, as shown in Fig. 1(b). However, it still employs transformer to perform cross-attention, owing to mismatched query and feature lengths-therefore it still incurs the quadratic cost of softmax attention. Additionally, the aforementioned approaches rely solely on the camera and LiDAR features of the most recent frame, neglecting valuable temporal sensor information from preceding frames. In dynamic scenarios—such as a pedestrian stepping out from behind a parked truck or a cyclist accelerating through an intersection—single-frame sensor information can fail to anticipate motion patterns, leading to unsafe or overly conservative plans. Integrating multi-frame history information is therefore essential to robustly predict scene evolution and ensure planner reliability. Moreover, when a transformer-based backbone leverages multi-frame camera and LiDAR features to enrich BEV features, its quadratic attention cost incurs prohibitively high computation, further compromising real-time performance. Thus, it is crucial to develop an end-to-end framework that efficiently fuses multi-frame sensor information, especially for resource-limited edge devices.
In recent developments of recurrent-style language models [gu2023mamba, dao2024transformers, peng2023rwkv, peng2024eagle, peng2025rwkv, kimiteam2025kda], linear attention has gained widespread attention for achieving transformer‑style parallel efficiency during training while preserving the lightweight, efficient inference characteristic of RNNs. In this work, we aim to leverage the latest RWKV-7 modules to fuse multi-frame camera and LiDAR features, enhancing planning performance while substantially reducing computational overhead. Consequently, we propose LADY, a fully linear attention-based framework that fuses multi-frame camera and LiDAR features to generate multi-modal trajectories, as shown in Fig. 1(c). Notably, our framework is designed with good extensibility and can be seamlessly adapted to other linear attention models beyond RWKV-7, offering flexibility for further performance optimization and model diversification.
Our main contributions are as follows:
  • To our best knowledge, LADY is the first fully linear attention-based end-to-end autonomous driving model.
  • We propose a lightweight linear cross-attention (LICA) mechanism to guarantee LADY’s end-to-end linear attention, and integrate it with a diffusion-based decoder to enhance the multi-modal distribution of trajectories.
  • LADY is designed to support fusion of long‑range temporal context at inference with constant computational and memory overhead—regardless of the length of historical camera and LiDAR features—thereby enhancing planning performance.
  • Experimental results demonstrate that LADY achieves notable performance improvement in both planning accuracy and runtime efficiency on the NAVSIM and Bench2Drive benchmarks. Additionally, LADY is validated on edge devices, confirming its applicability in the real world.
II Related Works
II-A End-to-End Autonomous Driving
By leveraging advances in perception and computational resources, end-to-end autonomous driving approaches can directly map raw sensor data to feasible trajectories. Following the taxonomy of [jiang2025transdiffuser], current approaches fall into three main categories: regression-based approaches, scoring-based approaches, and diffusion-based approaches.
Regression-based approaches [chitta2022transfuser, hu2023planning, ye2023fusionad, jiang2023vad, yuan2024drama] predict a trajectory in an auto-regressive fashion. UniAD [hu2023planning] integrates multiple perception tasks—such as object detection and semantic segmentation—into a unified planning network, demonstrating the great potential of end-to-end autonomous driving. Transfuser [chitta2022transfuser] fuses camera and LiDAR features at multiple scales to generate a feasible trajectory via an auto-regressive paradigm and has become a standard baseline on the NAVSIM benchmark [dauner2024navsim], which aligns open-loop and closed-loop evaluation metrics and provides a public evaluation server for end-to-end planning.
To capture diverse driving behaviors, scoring-based approaches [li2024hydra, chen2024vadv2, li2025hydra, guo2025ipad] generate multi-modal candidate trajectories and select the optimal trajectory using a learned or hand-crafted scoring function. VADv2 [chen2024vadv2] extends VAD [jiang2023vad] by generating multi-modal trajectories: it scores and samples from a large, fixed vocabulary of anchor paths to improve planning performance. Moreover, the Hydra-MDP series [li2024hydra, li2025hydra] further refines the scoring mechanism by incorporating extra supervision from a rule-based planner, enabling the model to learn from both expert heuristics and human driving behavior in a scalable manner. Besides, iPad [guo2025ipad] introduces a dedicated scorer to assess the safety, efficiency, and comfort of each predicted proposal, with ground-truth scores computed following NAVSIM [dauner2024navsim].
Inspired by generative diffusion models, diffusion-based approaches [liao2024diffusiondrive, jiang2025transdiffuser] directly generate multi-mode trajectories from a Gaussian distribution via iterative denoising process. DiffusionDrive [liao2024diffusiondrive] first introduces the diffusion model to the field of end-to-end autonomous driving and proposes a novel truncated diffusion policy to mitigate mode collapse and reduce inference cost. Additionally, TransDiffuser [jiang2025transdiffuser] introduces a simple yet effective multi-modal representation decorrelation optimization mechanism during the training process to further improve the diversity and quality of generated trajectories.
II-B Application of Linear Attention in Autonomous Driving
Figure 2: Overall architecture of our approach LADY. The model’s inputs comprise the ego vehicle’s current status, consecutive multi-frame camera and LiDAR features, and noisy anchor trajectories; its outputs include denoised multi-modal trajectories, passability maps and agent future states related to potential collisions. During training, multi-frame camera and LiDAR features are fused in parallel. During inference, the current frame camera and LiDAR features are fused sequentially with temporal hidden state to generate the BEV feature.
Despite significant progress in generating multi-modal candidate trajectories, existing approaches still rely on transformer-based feature extraction and cross-attention, whose quadratic complexity undermines real-time performance. Linear attention architectures include RWKV [peng2024eagle, peng2025rwkv], RetNet [sun2023retentive], Gated Linear Attention [yang2024gated], Mamba-2 [dao2024transformers], and KDA [kimiteam2025kda] remedy this by reformulating attention as an RNN with constant running time per token and constant memory usage.
DRAMA [yuan2024drama] employs Mamba-2 to effectively fuse camera and LiDAR features, demonstrating improved computational efficiency. However, Mamba only focuses on self-attention within each individual modality. For those multi-modal which requires cross-attention, which are common in end‑to‑end autonomous driving, DRAMA still relies on conventional transformer‑based cross‑attention. Consequently, it cannot be regarded as a fully linear attention-based end-to-end framework. Furthermore, DRAMA’s regression-based design inherently limits its performance, preventing it from achieving state-of-the-art (SOTA) results.
Additionally, while Mamba-2 improves computational efficiency, its per time-step decay scheme merely suppresses older state contents instead of removing them entirely, resulting them becoming a smaller proportion of the numerically increasing state. DeltaNet [schlag2021linear] addresses this by applying the Error Correcting Delta Rule [widrow1988adaptive] to compress key–value states, allowing both addition and removal of information on a per-key basis. Moreover, parallelizing DeltaNet [yang2024parallelizing] employs a diagonal plus low-rank state update, enabling efficient training of these models. Building on these insights, RWKV-7 [peng2025rwkv] further extends that parallelization to cover the generalized delta rule formulation and overcomes Mamba-2’s retention issue while preserving linear complexity.
Besides, existing approaches typically process only a single recent frame of camera and LiDAR features. As a result, they lack temporal context that is crucial for predicting dynamic agents and anticipating occluded hazards. To improve planning performance while further enhancing computational efficiency, we employ the latest RWKV-7 model to fuse multi-frame camera and LiDAR features. Moreover, to maintain end-to-end linear attention, we introduce a lightweight LICA mechanism with linear time and memory complexity to align modalities efficiently. Therefore, LADY adopts a fully linear attention architecture.
III Method Design
The overall architecture of LADY is generally an encoder-decoder model, as shown in Fig. 2. The encoder of LADY fuses raw multi-frame camera and LiDAR features to generate BEV features, while the decoder uses these features to predict multi-modal trajectories. Specifically, during training, LADY fuses multi-frame camera and LiDAR features in parallel to improve efficiency, as demonstrated in Fig. 2(a). In contrast, at inference time, LADY fuses the current frame camera and LiDAR features with the temporal hidden state sequentially, as shown in Fig. 2(b). Therefore, LADY is able to fuse long-range temporal context at inference. Below, we provide a detailed explanation of the key components in the architecture of LADY.
III-A Linear Attention-based Multi-Frame Feature Fusion
RWKV Mechanism To efficiently fuse multi-frame camera and LiDAR features, we employ the lightweight, RNN-style RWKV [peng2024eagle, peng2025rwkv] instead of standard transformers [vaswani2017attention]. RWKV model stacks multiple residual blocks, each integrating a time mixing sub-block and a channel mixing sub-block to form an RNN-like, time sequential architecture that effectively leverages past information, as shown in Fig. 3.
Figure 3: Overall architecture of the RWKV-7 block.
Recently, RWKV-7 (“Goose”) extends the basic RWKV architecture by introducing the delta rule for key–value compression. We first introduce two common operators used in RWKV-7:
$\text{lerp}(\bm{a},\bm{b},\bm{\mu})$ $=\bm{a}+(\bm{b}-\bm{a})\odot\bm{\mu},$
$\text{loramlp}_{\square}(f,\bm{x},\textbf{bias})$ $=f(\bm{x}\bm{A}_{\square})\bm{B}_{\square}+(\bm{\lambda}\;\text{if {bias} else}\;\bm{0}),$
where $\odot$ denotes element-wise multiplication, $\text{lerp}(\bm{a},\bm{b},\bm{\mu})$ denotes the linear interpolation (lerp) between $\bm{a}$ and $\bm{b}$, and here is used to implement the token shift operation; $\text{loramlp}_{\square}(f,\bm{x},\textbf{bias})$ represents the low-rank MLP, and $\square$ represents the elements of the RWKV model. At each time step $t$, the RWKV model is typically defined by seven fundamental elements: the receptance vector $\bm{r}_{t}$, the weight decay vector $\bm{w}_{t}$, the replacement key vector $\tilde{\bm{k}}_{t}$, the removal key vector $\bm{\kappa}_{t}$, the value vector $\bm{v}_{t}$, the in-context learning rate vector $\bm{a}_{t}$, and the gate vector $\bm{g}_{t}$. Additionally, each element can be computed as follows:
$\bm{x}_{t}^{\square}$ $=\text{lerp}(\bm{x}_{t},\bm{x}_{t-1},\bm{\mu}_{\square}),\;\square\in\{r,w,k,v,a,g\},$
$\bm{r}_{t}$ $=\bm{x}_{t}^{r}\bm{W}_{r},$
$\bm{w}_{t}$ $=\text{exp}(-e^{-0.5}\text{sigmoid}(\text{loramlp}_{w}(\text{tanh},\bm{x}_{t}^{w},\textbf{bias}=\text{True}))),$
$\bm{k}_{t}$ $=\bm{x}_{t}^{k}\bm{W}_{k},$
$\bm{\kappa}_{t}$ $=\bm{k}_{t}\odot\bm{\xi},$
$\bm{a}_{t}$ $=\text{sigmoid}(\text{loramlp}_{a}(\text{Identity},\bm{x}_{t}^{a},\textbf{bias}=\text{True})),$
$\tilde{\bm{k}}_{t}$ $=\bm{k}_{t}\odot\text{lerp}(\bm{1},\bm{a}_{t},\bm{\alpha}),$
$\bm{\nu}_{t}$ $=\text{sigmoid}(\text{loramlp}_{v}(\text{Identity},\bm{x}_{t}^{v},\textbf{bias}=\text{True}),$
$\bm{v}_{t,l}$ $=\bm{x}_{t}^{v}\bm{W}_{v},$
[…]
$\bm{g}_{t}$ $=\text{loramlp}_{g}(\text{sigmoid},\bm{x}_{t}^{g},\textbf{bias}=\text{False}),$
where $\bm{\xi}$ and $\bm{\alpha}$ are learnable parameters. The temporal hidden state $\bm{S}_{t}$ is then obtained through the following sequential form:
$\bm{S}_{0}=\bm{0},\,\bm{S}_{t}=\bm{S}_{t-1}(\text{diag}(\bm{w}_{t})-\hat{\bm{\kappa}}_{t}^{\top}(\bm{a}_{t}\odot\hat{\bm{\kappa}}_{t}))+\bm{v}_{t}^{\top}\cdot\tilde{\bm{k}}_{t},$ (1)
where $\hat{\bm{\kappa}}_{t}=\bm{\kappa}_{t}/\|\bm{\kappa}_{t}\|_{2}$. Additionally, the entire state update process could be fully computed in a parallel form, which could use Tensor Cores on GPU to accelerate. Let the chunk span local indices $i\in\{0,\dots,B-1\}$, and we define the causal mask decay matrix as:
$\Delta_{ij}=\prod_{m=j}^{i}\bm{w}_{m}\>(j\leq i),\>\Delta_{ij}=0\>(j>i).$ (2)
Besides, let $\bm{u}_{t}=\bm{a}_{t}/\|\bm{\kappa}_{t}\|_{2}$, and we can obtain the stacked matrices $\mathbf{U},\tilde{\mathbf{K}},\mathbf{V}\in\mathbb{R}^{B\times d}$. Given the incoming temporal hidden state $\mathbf{S}_{\text{in}}\in\mathbb{R}^{d\times d}$, the parallel form for the entire chunk is written as:
$\mathbf{S}=\Delta\odot\bigl[\mathbf{S}_{\text{in}}+\tilde{\mathbf{K}}^{\top}\mathbf{V}-\mathbf{S}_{\text{in}}(\mathbf{U}^{\top}\mathbf{V})\bigr]\bigl(\mathbf{I}+\mathbf{V}^{\top}(\Delta\odot\mathbf{U})\bigr)^{-1}.$
The rows of $\mathbf{S}\in\mathbb{R}^{B\times d}$ are the hidden states $\mathbf{S}_{0},\dots,\mathbf{S}_{B-1}$ that the recurrent rule would have produced step-by-step. For further details regarding its implementation in PyTorch, please refer to FLA [yang2024fla]. The output of time mixing is then computed as follows:
$\bm{p}_{t}$ $=\text{LayerNorm}(\bm{r}_{t}\bm{S}_{t}^{\top})+(\bm{r}_{t}\cdot(\bm{\rho}\odot\tilde{\bm{k}}_{t})^{\top})\bm{v}_{t},$ (3)
$\bm{o}_{t}$ $=(\bm{g}_{t}\odot\bm{p}_{t})\bm{W}_{o},$
where $\bm{\rho}$ is also a learned parameter. Moreover, the output of the time mixing sub-block is then passed through the channel mixing sub-block to obtain the final RWKV block output:
$\bm{o}^{\prime}_{t}=\text{ReLU}(\text{lerp}(\bm{x}^{\prime}_{t},\,\bm{x}^{\prime}_{t-1},\,\bm{\mu}_{k^{\prime}})\bm{W}_{k^{\prime}})^{2}\mathbf{W}_{v^{\prime}}.$ (4)
The complete RWKV block architecture is illustrated in Fig. 3, and for further details of the specific meaning of each element, please refer to [peng2025rwkv].
Multi-Frame Feature Fusion We adopt the same ResNet-34 [he2016deep] backbone as Transfuser [chitta2022transfuser] to extract features from multi-frame camera and LiDAR features. Additionally, we employ RWKV-7 to fuse multi-frame camera and LiDAR features with linear time complexity, as shown in Fig. 2, enabling comprehensive extraction of both spatial and temporal information across consecutive frames. During the training stage, we organize the camera and LiDAR features into separate sequences across different time steps. Then, for each time step, we concatenate the corresponding camera and LiDAR features to form an augmented representation, as shown in Fig. 2(a). Besides, we add learnable spatial positional embeddings to the concatenated multi-frame features. The concatenated multi-frame features are then fed into stacked RWKV-7 blocks to efficiently extract information across both temporal and cross-modal dimensions. After fusion, we split the resulting sequence back into individual camera and LiDAR streams for each frame, as illustrated in Fig. 2(a). Compared to the training stage, where multi-frame camera and LiDAR features are fused in parallel, the inference stage of LADY performs sequential fusion—integrating current camera and LiDAR features conditioned on temporal hidden state, as illustrated in Fig. 2(b). Thus, LADY facilitates the integration of long-range temporal context, which is critical for autonomous driving.
After performing dense feature fusion at multiple resolutions, we apply convolution to the fused LiDAR feature to obtain the BEV feature map, as shown in Fig. 2(a). We then concatenate the BEV map with the current ego status, and add learnable positional embeddings to the augmented representation, providing richer context for downstream processing. Finally, to enhance robustness, we adopt the feature state dropout (FSD) strategy [yuan2024drama], randomly dropping either BEV or ego-state channels with differentiated probabilities.
III-B Linear Cross-Attention Mechanism
Figure 4: Overall architecture of linear cross-attention block.
The decoder is responsible for generating multi-model trajectory candidates, each predicted trajectory is represented as a sequence of $N$ waypoints $\tau=\{(x_{t},y_{t},\theta_{t})\}_{t=1}^{N}$, where $(x_{t},y_{t})$ is the position of each waypoint at time step $t$ in its own ego-centric coordinate frame, $\theta_{t}$ represents the orientation, and $N$ denotes the prediction horizon. First, we employ a learnable query to perform cross-attention over the augmented BEV map and ego status representations, as shown in Fig. 2(a), obtaining queries that capture the features related to surrounding agents. These queries then serve as conditions for the diffusion-based decoder to generate multi-modal trajectories. As the query sequence length differs from the augmented feature sequence, existing approaches [yuan2024drama, dao2024transformers] typically employ transformer to perform cross-attention [vaswani2017attention]. However, employing transformer-based cross-attention would violate LADY’s fully linear-attention design, making it difficult to ensure real-time application when fusing with multi-frame BEV features. To further improve efficiency while ensuring fully linear-attention design, we propose LICA, as shown in Fig. 4, which can also be extended to all linear attention architectures.
The query is first passed through an RWKV-7 block to encode its positional dependencies—analogous to self-attention in transformers—producing encoded query. Additionally, we concatenate the fused BEV-ego representation with the encoded query along the sequence dimension, and then the combined sequence is fed through a second RWKV-7 block, whose recurrent state updates naturally attend to both earlier feature tokens and later query tokens, thus implementing LICA without quadratic cost. We then extract the last $M$ tokens of the output sequence to serve as the cross-attended queries, where $M$ equals the original query length, as shown in Fig. 4. Therefore, our approach achieves efficient LICA that scales linearly with sequence length, enabling the fully linear attention of LADY.
III-C Diffusion-based Trajectory Decoder
Figure 5: Overall architecture of the diffusion-based decoder layer.
To capture richer multi-modal trajectory distributions, we apply the truncated diffusion policy from DiffusionDrive [liao2024diffusiondrive]. The diffusion-based decoder takes slightly noisy anchor trajectories as the input, and at each denoising step, conditions on the fused BEV features and the query related to surrounding agents. Specially, we leverage our proposed LICA block to perform cross-attention between the noisy anchor trajectories and both the BEV feature and agent queries, as shown in Fig. 5. Additionally, inspired by iPad [guo2025ipad], the diffusion-based decoder outputs a set of denoised multi-modal trajectories along with their confidence scores, and two optional output heads, as shown in Fig. 2. Scores are distilled from rule-based driving knowledge embedded in NAVSIM [dauner2024navsim], and the trajectory with the highest score is selected as the final predicted trajectory. Besides, as shown in Fig. 2, the mapping head is used to predict the on-road and on-route probabilities for multi-modal trajectories. While the prediction head is used to predict agents’ future states related to potential collisions, therefore enabling more aggressive, yet safe, ego-vehicle maneuvers.
IV Experiments
TABLE I: Comparison on NAVSIM navtest split with the start-of-the-art. C & L represents camera and LiDAR information. Multi-frame means that the model use multi-frame sensor information to generate feasible trajectories. Besides, $T$ denotes the sequence length and $d$ is the feature dimension.
Methods Modality Architecture BackBone Multi-Frame Complexity Metrics
Time Memory NC DAC TTC Comf. EP PDMS
UniAD [hu2023planning] Camera Transformer ResNet-34 $\mathcal{O}(T^{2}d)$ $\mathcal{O}(T^{2}+Td)$ 97.8 91.9 92.9 100 78.8 83.4
Transfuser [chitta2022transfuser] C & L Transformer ResNet-34 $\mathcal{O}(T^{2}d)$ $\mathcal{O}(T^{2}+Td)$ 97.7 92.8 92.8 100 79.2 84.0
DRAMA [yuan2024drama] C & L MTb ResNet-34 $\mathcal{O}(Td)$c $\mathcal{O}(d)$ 98.0 93.1 94.8 100 80.1 85.5
VADv2 [chen2024vadv2] Camera Transformer ResNet-34 $\mathcal{O}(T^{2}d)$ $\mathcal{O}(T^{2}+Td)$ 97.2 89.1 91.6 100 76.1 80.9
Hydra-MDP [li2025hydra] C & L Transformer ResNet-34 $\mathcal{O}(T^{2}d)$ $\mathcal{O}(T^{2}+Td)$ 98.3 96.0 94.6 100 78.7 86.5
DiffusionDrive [liao2024diffusiondrive] C & L Transformer ResNet-34 $\mathcal{O}(T^{2}d)$ $\mathcal{O}(T^{2}+Td)$ 98.2 96.2 94.7 100 82.2 88.1
iPad [guo2025ipad] Camera Transformer ResNet-34 $\mathcal{O}(T^{2}d)$ $\mathcal{O}(T^{2}+Td)$ 98.6 98.3 94.9 100 88.0 91.7
DiffusionDrivea C & L Transformer ResNet-34 $\mathcal{O}(T^{2}d)$ $\mathcal{O}(T^{2}+Td)$ 97.5 97.1 93.2 100 86.1 89.3
LADY (Ours) C & L Linear Attention ResNet-34 $\mathcal{O}(Td)$ $\mathcal{O}(d)$ 98.0 97.3 94.0 100 88.6 90.9
  • a
    Multi-frame version   $^{\text{b}}$ Mamba & Transformer
  • c
    DRAMA’s Transformer-based cross-attention makes its time and memory complexity slightly higher than the table values and LADY.
IV-A Experiment Setup
Dataset We evaluate LADY on the NAVSIM benchmark [dauner2024navsim]. NAVSIM is built upon OpenScene, a compact redistribution of the public autonomous driving dataset nuPlan, which provides 120 hours of driving logs, recorded at a lower frequency of 2Hz. The input of the ResNet-34 backbone includes multi-frame camera and LiDAR features, and each frame contains three cropped and downscaled forward-facing camera images, concatenated as a 1024×256 image, and a rasterized LiDAR. Furthermore, the input also incorporates the current ego vehicle’s status, including velocity, acceleration, and driving commands from the navigation module, such as turning, lane changing and following.
Metrics NAVSIM provides the Predictive Driver Model Score (PDMS) to evaluate the model’s performance, which is defined as:
$\mathrm{PDMS}=\underbrace{\left(\prod_{m\in\{\text{NC},\text{DAC}\}}\mathsf{score}_{m}\right)}_{\text{penalties}}\times\underbrace{\left(\frac{\sum_{w\in\{\text{EP},\text{TTC},\text{C}\}}\mathsf{weight}_{w}\times\mathsf{score}_{w}}{\sum_{w\in\{\text{EP},\text{TTC},\text{C}\}}\mathsf{weight}_{w}}\right)}_{\text{weighted average}},$
where driving with no collision (NC) with traffic participants and no infraction regarding drivable area compliance (DAC) are two hard penalties. If either requirement is violated, $\mathsf{score}_{\text{NC}}$ or $\mathsf{score}_{\text{DAC}}$ is set to 0, which in turn yields a PDMS of 0. Additionally, three sub-scores capture specific aspects of the planned trajectory: ego progress (EP) assesses the ego progress along the route center; time-to-collision (TTC) measures the minimum safety margin relative to other vehicles; comfort (C) evaluates the ride quality by comparing the acceleration and jerk of the trajectory to predetermined thresholds.
IV-B Implementation Details
The input of LADY consists of multi-frame sensor information alongside noisy anchor trajectories. For the final evaluation, the output is an 8-waypoint trajectory spanning $4$ seconds, sampled at 2Hz, with each waypoint defined by x, y and heading coordinates. Additionally, LADY is trained from scratch for 100 epochs on the navtrain split using the Adam optimizer across $4$ NVIDIA RTX 4090 GPUs ($48$ GB each), with a total batch size of $4$. We set the initial learning rate to $1\times 10^{-4}$ and employ a ReduceLROnPlateau scheduler ($\texttt{patience}=10$ and $\texttt{factor}=0.5$). Specially, LADY is trained with a fixed history length of $10$ frames—comprising the current frame and the preceding $9$ time steps frames—covering a $5.0$-second span at 2Hz; data with fewer than 10 frames is prepended with zeros to meet the required length. Furthermore, for the diffusion-based trajectory decoder, we stack $2$ cascaded diffusion decoder layers and employ a truncated diffusion policy as DiffusionDrive [liao2024diffusiondrive]: during training, we truncate the diffusion schedule to $50$ out of $1000$ steps to diffuse the $100$ clustered anchor trajectories, thereby enriching multi-modal representations. At inference, we only perform 2 denoising steps and select the highest-scoring predicted trajectory for evaluation. Evaluation is conducted on the NVIDIA Jetson AGX Orin ($32$ GB), a widely adopted embedded platform for autonomous driving applications. Finally, LADY integrates the RWKV architecture following Flash Linear Attention [yang2024fla].
IV-C Open-Loop Quantitative Results
Figure 6: Inference time and memory usage comparison between Transfuser, DRAMA, DiffusionDrive, and our approach LADY.
Table I compares LADY with SOTA methods on the NAVSIM navtest split for open-loop evaluations. All models use the same ResNet-34 backbone, and LADY achieves an overall PDMS of $90.9$. By leveraging multi-frame sensor information and a diffusion-based decoder, LADY outperforms most current SOTA methods. Additionally, compared to Transfuser and DRAMA, the significant improvement in the EP metric demonstrates that our method is considerably less conservative. Besides, Fig. 6 further compares the inference time and memory usage for LADY, Transfuser, DRAMA and DiffusionDrive as input frame count increases, with all measurements conducted on the NVIDIA Orin platform. During inference, i.e., in the sequential mode, LADY maintains a compact temporal hidden state that is updated frame-by-frame. Therefore, its computational graph is independent of the total number of frames for both camera and LiDAR features, resulting in virtually constant latency and memory footprint as the frame count increases. In contrast, the transformer-based architectures of Transfuser and DiffusionDrive require recomputing attention over all frame tokens, causing theirs inference time and memory consumption to grow sharply with each additional frame. For DRAMA, as it does not adopt a full linear attention architecture, its computational overhead and memory usage also rise with increasing input frames. Therefore, the constant-time, constant-memory profile makes LADY especially well suited for real-time, end-to-end autonomous driving, where fusing long-range temporal context is critical.
IV-D Closed-Loop Bench2Drive Results
TABLE II: Closed-loop and multi-ability results of different E2E methods on Bench2Drive benchmark.
Method Closed-loop Results Multi-Ability Results (%)
[…]
AD-MLP 22.63 48.45 0.00 18.05 0.00 0.00 0.00 0.00 4.35 0.87
UniAD-Tiny 47.04 123.92 13.18 40.73 8.89 9.33 20.00 20.00 15.43 14.73
UniAD-Base 43.58 129.21 16.36 45.81 14.10 17.78 21.67 10.00 14.21 15.55
VAD 46.01 157.94 15.00 42.35 8.11 24.44 18.64 20.00 19.15 18.07
DriveTransformer 20.78 100.64 35.01 63.46 17.57 35.00 48.36 40.00 52.10 38.60
iPad 28.21 161.31 35.91 65.02 30.00 20.00 53.33 60.00 49.47 42.56
LADY(1 frame) 17.55 141.31 14.55 50.06 20.00 4.44 8.33 50.00 24.21 21.40
LADY(10 frames) 36.26 155.39 25.91 57.52 26.25 33.33 21.67 50.00 28.42 31.93
LADY(Infinite frames) 12.13 155.57 33.02 65.12 24.36 55.81 35.71 70.00 28.80 42.94
TCP* 47.80 54.26 15.00 40.70 16.18 20.00 20.00 10.00 6.99 14.63
TCP-ctrl* 51.51 55.97 7.27 30.47 10.29 4.44 10.00 10.00 6.45 8.23
TCP-traj* 18.08 76.54 30.00 59.90 8.89 24.29 51.67 40.00 46.28 34.23
TCP-traj w/o distill. 22.96 78.78 20.45 49.30 13.53 16.67 40.00 50.00 28.72 28.51
ThinkTwice* 16.22 69.33 31.23 62.44 27.38 18.42 35.82 50.00 54.23 37.17
DriveAdapter* 16.01 70.22 33.08 64.22 28.82 26.38 48.76 50.00 56.43 42.08
    denotes expert feature distillation
Given the challenges associated with evaluating closed-loop driving performance in real-world scenarios, we instead leverage the CARLA simulator [dosovitskiy2017carla] and adopt the Bench2Drive benchmarks [jia2024bench2drive] for performance assessment. Consistent with the original work, we use the same training dataset and evaluation metrics. To systematically evaluate the impact of long-term temporal context utilization on performance, LADY is implemented in three variants with distinct temporal feature fusion strategies: 1 frame: only leverages the camera and LiDAR features extracted from the current frame (i.e., no historical information is incorporated); 10 frames: fuses features from the current frame and up to $9$ preceding frames, enabling short-term temporal context modeling; Infinite frames: maintains and updates a hidden state that accumulates feature information from all historical frames (in addition to the current frame), thus supporting long-term temporal context aggregation. As demonstrated in Table II, LADY benefits from its effective utilization of long-term temporal context, and consequently surpasses baseline methods. Notably, this advantage is achieved without introducing extra computational overhead, highlighting the efficiency of its temporal modeling design.
IV-E Ablation Study
TABLE III: Ablation study of fixed‑frame model with respect to varying frame number.
[…]
1 97.4 95.5 91.8 100 83.2 86.8
4 97.6 96.9 92.9 100 86.8 89.5
6 97.7 96.9 92.9 100 87.8 90.0
8 97.7 97.1 93.3 100 88.2 90.3
10 97.9 97.1 93.6 100 88.5 90.6
12 97.9 97.2 93.7 100 88.5 90.7
15 98.0 97.3 94.0 100 88.6 90.9
20 98.0 97.2 94.1 100 88.5 90.9
First, we train a multi-frame version of DiffusionDrive, achieving a score improvement from $88.1$ to $89.3$, validating the effectiveness of incorporating multi-frame information. Additionally, we evaluated the performance of LADY trained on a fixed history length of $10$ frames when inferring with varying input lengths, as shown in Table III. Specifically, if the frame count of the evaluation data fails to meet the requirement, zero-padding is still applied to the front for data filling. The results demonstrate that when the number of input frames at inference is less than the training length, performance degrades due to insufficient information. Conversely, when presented with more frames than seen during training, the model seamlessly incorporates the additional temporal hidden state with the current camera and LiDAR features, further improving planning performance. It even outperforms its baseline trained‑frame performance, demonstrating LADY’s suitability for real‑time inference in real‑world driving scenarios. Therefore, LADY can leverage long-range temporal context from multi-frame sensor information, a capability that transformer‑based approaches cannot match.
IV-F Qualitative Results
Figure 7: Visualization of representative NAVSIM scenarios of our proposed model LADY.
To qualitatively validate LADY, we provide the visualization of the multi-modal trajectories generated by LADY on $8$ representative scenarios drawn from the navtest split, as shown in Fig. 7. Across these scenarios, LADY generates multi-modal trajectories and selects the optimal trajectory based on confidence scores, thereby demonstrating safe and contextually appropriate behaviors—including following a leading vehicle, overtaking slower-moving vehicles, and executing turns. Furthermore, compared with the ground truth, LADY avoids excessive conservatism and significantly outperforms the current state of the art.
(a) Intersection scenario
(b) Long-horizon route scenario
Figure 8: Comparison of representative Bench2Drive scenarios between our proposed LADY model and iPad [guo2025ipad].
Additionally, we present two representative scenarios drawn from Bench2Drive, comparing our proposed approach with iPad [guo2025ipad], as illustrated in Fig. 8. For the intersection scenario, where only the camera and LiDAR features of the current frame are provided as input, iPad tends to adopt an overly conservative strategy, resulting in prolonged stationary behavior at the intersection. In contrast, our model leverages historical observation data to infer richer interaction contexts, thereby generating more reasonable and effective trajectory planning strategies. For the long-horizon route scenario, our method successfully completes the entire navigation task by accumulating historical observation sequences. By comparison, iPad exhibits hesitant behaviors at complex traffic-light intersections, leading to conservative decisions and eventual mission failure due to progress stagnation.
IV-G Limitations
Certainly, compared to SOTA methods, such as iPad [guo2025ipad], our performance still has a certain gap. However, when using the NAVSIM reference trajectory evaluator to select the optimal trajectory from multi-modal trajectory candidates, the PDMS achieves a score of $99.2$, as shown in Table IV. This result demonstrates that the current ResNet-34 backbone is able to generate high-quality candidate trajectories, yet the existing scorer fails to accurately identify the optimal one from them. Given that the core of this study focuses on the exploration of linear attention in the field of autonomous driving, the performance improvement of the scorer will be further carried out in future work. Additionally, it is worth noting that the modest performance gains are mainly attributed to the insufficient number of samples requiring long-range temporal context in the current dataset, we expect LADY’s advantages to become more pronounced in complex scenarios where temporal context is critical.
TABLE IV: Best trajectory in the multi-modal trajectory candidates.
Methods NC DAC TTC Comf. EP PDMS
LADY 98.0 97.3 94.0 100 88.6 90.9
LADY (Best-of-N) 100 100 100 100 98.13 99.2
Human 100 100 100 100 87.5 94.8
V Conclusion
In this work, we propose LADY, the first fully linear attention–based end-to-end autonomous driving model. Additionally, we introduce a lightweight LICA mechanism for linear cross-attention, preserving LADY’s end-to-end linearity. By leveraging linear attention, LADY fuses long-range camera and LiDAR features with constant computational and memory overhead. Experimental results on the NAVSIM planning and Bench2Drive benchmarks demonstrate that LADY achieves significantly lower inference cost on long-range sequences while delivering superior planning performance. Future work will focus on refining the scorer to improve performance, incorporating alternative linear‑attention architectures into our model, and validating LADY in real-world autonomous driving scenarios.